a metal piece of mass 10grm  is  suspended by a vertical string.the spring elongates 10cm over its natural length to keep the piece in equillibrium. a beaker containing water is now placed below the piece so as to immerse the piece completely in water. find the elongation of the spring. density of  metal =9000kg/m3.

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

volume of metal=10/9 cc

buoyant force on metal=(10/9)*1

for spring, F=kx,

tensile force on spring =gravity- buoyant force

=10(1-(1/9))*g=kx2

10*g=kx2, x2=10cm